OpenStudy (kpop4life123):

Find a rational number and an irrational number that are between 9.5 and 9.7. Include the decimal approximation of the irrational number to the nearest hundredth.

OpenStudy (anonymous):

9.6 is arational number \[\left( 9.6 \right)^2=92.16\] \[\sqrt{92.161}~is~ an~ irrational~ number\]